# TOMG-Bench: Evaluating LLMs on Text-based Open Molecule Generation

### v2

@@ Abstract
In this paper, we propose Text-based Open Molecule Generation Benchmark (TOMG-Bench), the first benchmark to evaluate the open-domain molecule generation capability of LLMs. TOMG-Bench encompasses a dataset of three major tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom). Each major task further contains three subtasks, while each subtask comprises 5,000 test samples. Given the inherent complexity of open molecule generation evaluation, we also developed an automated evaluation system that helps measure both the quality and the accuracy of the generated molecules. Our comprehensive benchmarking of 25 LLMs reveals the current limitations as well as potential areas for improvement in text-guided molecule discovery. Furthermore, we propose OpenMolIns, a specialized instruction tuning dataset established for solving challenges raised by TOMG-Bench. Fine-tuned on OpenMolIns, Llama3.1-8B could outperform all the open-source general LLMs, even surpassing GPT-3.5-turbo by 46.5% on TOMG-Bench. Our codes and datasets are available through https://github.com/phenixace/TOMG-Bench.
@@ 1 Introduction
Molecule discovery plays a pivotal role in various scientific research fields, from pharmaceuticals (Keiser et al., 2010) to materials science (Higuchi et al., 2023). Normally, molecule discovery is a trial-and-error process (Ekins, 2024), which requires repetitive experimentation and data analysis (Mattern & Grosser, 2023). Due to the inefficiency of the traditional techniques, it usually takes more than 10 years to bring a new drug candidate into the market (Lee et al., 2018).
With the development of machine learning techniques and the advent of Graph Neural Networks (GNNs) (Wang et al., 2023), there has been a significant step forward. As molecules could be represented as graphs, GNN-based methods can capture the structural patterns of the molecule and make accurate predictions. With the assistance of GNNs, researchers could analyse the properties of molecules (Cai et al., 2022) and generate new molecule candidates (Jin et al., 2018). However, challenges still exist. GNN-based methods struggle to generalize to different tasks (Chen et al., 2024), necessitating costly data collection and preparation for different downstream tasks. Moreover, these methods are constrained in their capacity to generate molecules with specific, customized properties, limiting their flexibility in molecular design (Li et al., 2024c).
In contrast, Large Language Models (LLMs) have shown their great generalization capability (OpenAI, 2023) and could be easily adapted to different research fields. For instance, Cascella et al. (2023) utilize ChatGPT for supporting clinical practice and Zhen et al. (2024) adopts LLMs as assistants for task planning in the field of Civil Engineering, showing the great potential of LLMs in scientific discovery.
Figure 1: Comparison of Text-based Targeted Molecule Generation (a) v.s. Text-based Open Molecule Generation (b).
As molecules can be represented as texts by Simplified Molecular Input Line Entry System (SMILES), a linear notation that encapsulates the structure of a chemical compound (Weininger, 1988), they can be processed and understood by LLMs, bridging the gap between molecules and natural languages. With advanced reasoning and in-context learning capabilities, LLMs are particularly adept at generalizing to the molecule domain (Li et al., 2024b). The generalization capability makes LLMs a viable option in molecule discovery. Furthermore, by aligning molecules with textual data, LLMs can serve as powerful assistants to chemists (Zhang et al., 2024; Li et al., 2024d). They could help interpret and generate chemical knowledge, suggest modifications to molecular structures, and even predict the properties and behaviours of compounds, which would potentially streamline the molecule discovery process, leading to breakthroughs in diverse research areas.
While the integration of LLMs into molecule discovery holds immense promise, the process of aligning molecules with textual data is challenging (Li et al., 2024a). A significant challenge lies in the availability and diversity of datasets and benchmarks necessary for training and evaluation. Although the molecule-caption translation datasets including ChEBI-20 (Edwards et al., 2022), PubChem324K (Liu et al., 2023) and L+M-24 (Edwards et al., 2024) are crucial for bridging the gap between the molecular and textual domains, they still face several limitations that need to be addressed:
On the one hand, there is a concern about the generalization of models trained on the molecule-caption translation task (Li et al., 2024b). In real-world scenarios, molecule captions that describe molecular structures can be highly ambiguous, with multiple correct interpretations, while the current molecule-caption translation is actually a targeted generation task. In this case, these models often struggle to generalize to customized molecules, even for seemingly simple examples (Li et al., 2024b). This suggests a fundamental mismatch between the molecular and textual spaces, raising questions about whether this task could truly guide LLMs well. On the other hand, a critical issue is the inability to propose new molecule structures. The ultimate goal of molecule discovery is not just to understand and describe existing chemical compounds but to innovate and discover new ones, particularly in the context of drug discovery, indicating that the current molecule-caption translation task and the corresponding evaluation metrics fall short in this regard. Therefore, addressing these challenges is essential for harnessing the full potential of LLMs in molecule discovery.
In our efforts to bridge the gap between the natural language and the molecular spaces and to further facilitate LLMs as chemist assistants in molecule discovery, we propose a novel benchmark, Text-based Open Molecule Generation Benchmark (TOMG-Bench). TOMG-Bench is designed to evaluate the open-domain generative capabilities of LLMs in the molecular domain through a series of structured instructions for molecule design and operations. As shown in Figure 1, different from the previous targeted molecule generation tasks, Text-based Open Molecule Generation does not set a specific target or enables LLMs to generate an exactly matched molecule. Instead, we design automatic evaluation processes with chemical toolboxes like RDKit (Landrum, 2013) to test whether the generation meets the requirements. In other words, there can be multiple correct answers for a single question, and LLMs are only required to generate one of them. Notably, TOMG-Bench is meticulously categorized into three primary tasks, i.e., molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom). Each category contains three subtasks, and each subtask is composed of 5,000 test samples, providing a comprehensive assessment of whether LLMs truly grasp the molecular space. Meanwhile, we also propose a different set of evaluation metrics to evaluate and rank the performance of LLMs, which considers both the accuracy and quality of the generated molecules. To solve the challenges raised by TOMG-Bench, we propose an instruction-tuning dataset, OpenMolIns, by extracting and reformatting molecules with meticulously designed rules from an existing molecule database. OpenMolIns is structured across five distinct data levels (i.e., light, small, medium, large, and xlarge) to study the data scaling law of LLMs in the text-based open molecule generation task.
Our contributions are primarily threefold:
@@
    TOMG-Bench introduces the task of Text-based Open Molecule Generation, which consists of three major tasks and their nine subtasks in total to synthetically benchmark the performance of LLMs in this task, demonstrating the limitations of the targeted molecule generation datasets like ChEBI-20.
@@
    We benchmarked a suite of 25 LLMs, providing a brand-new perspective for the evaluation of LLMs.
@@
    We introduce OpenMolIns, an instruction tuning dataset, enabling Llama-3.1-8B-Instruct to outperform all the open-source general LLMs in TOMG-Bench.
2 Related Work
2.1 Development of AIs in Molecule Discovery
Molecule discovery plays a pivotal role across numerous scientific fields, driving advancements in the development of drug discovery and material design (Du et al., 2022). Thus, integrating artificial intelligence into molecule discovery has marked a transformative shift in the pharmaceutical landscape, significantly enhancing the efficiency and effectiveness of identifying and developing new therapeutic molecules. Recent advancements in machine learning (ML), deep learning (DL), and natural language processing (NLP) have enabled AI systems to analyze complex biological and chemical data more effectively than traditional methods (Wigh et al., 2022; Wu et al., 2018; Zhou et al., 2023). For instance, MolReGPT  (Li et al., 2024b) leverages LLMs like ChatGPT to learn molecule SMILES strings representation for molecule-caption translation tasks. Moreover, existing studies have explored advanced methods that utilize various AI techniques to further enhance molecule discovery processes, including Convolutional Neural Networks (CNNs) (Peng & Zhao, 2019; Le et al., 2019), Recurrent Neural Networks (RNNs) (Grisoni et al., 2020; Popova et al., 2019), Graph Neural Networks (GNNs) (Wang et al., 2023; Sun et al., 2022), and Transformer-based networks (Xia et al., 2023; Balaji et al., 2023; Edwards et al., 2022).
2.2 Text-based Molecule Generation
Text-based Molecule Generation (Edwards et al., 2021) has recently emerged as a transformative task to molecule discovery. Previously, this task centres on generating targeted molecules based on natural language descriptions, requiring the creation of paired datasets of molecules and their corresponding textual representations. Early approaches leveraged transformer-based models like MolT5 (Edwards et al., 2022), employing self-supervised learning on large datasets to generate high-quality SMILES strings from textual inputs. Subsequent advancements, such as KV-PLM (Zeng et al., 2022), MoMu (Su et al., 2022), and BioT5 (Pei et al., 2023), integrated molecular graphs or biochemical text to improve both understanding and generation capabilities. 3D-MoLM (Li et al., 2024e) further enhanced this by incorporating spatial configurations, leading to more accurate and geometrically valid molecular representations. Meanwhile, the application of LLMs like MolReGPT (Li et al., 2024b) and ICMA (Li et al., 2024a) as in-context learners has also shown significant promise. These models demonstrate the ability of LLMs to adaptively generate molecules by retrieving and leveraging relevant examples from the provided context. Most recently, MolReFlect (Li et al., 2024c) underscored the importance of fine-grained alignment between molecular structures and their textual descriptions, utilizing a teacher-student training paradigm to capture these nuanced relationships effectively.
3 TOMG-Bench
In this section, we propose TOMG-Bench to comprehensively assess the performance of LLMs in molecular space. Specifically, the benchmark is composed of three basic tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom). To ensure the integrity and effectiveness of TOMG-Bench, we have developed a robust set of evaluation metrics for different tasks. Additionally, we created OpenMolIns, an instruction-tuning dataset aimed at enhancing the performance and adaptability of LLMs to the challenges presented by this benchmark.
Figure 2: Task illustrations of TOMG-Bench. Both molecules to the right of the arrow are acceptable solutions.
3.1 Dataset Categorization
The categorization of the TOMG-Bench dataset initially considers the inherent characteristics of molecule SMILES representation and the role of LLMs serving as the chemist’s assistant, namely helping the chemists to edit, optimize and customize molecules as they want. As shown in Figure 2, following the difficulty of the tasks, we demonstrate the three basic tasks as well as their corresponding subtasks:
MolEdit emerges as the most straightforward task among the three domains, as an existing molecule is already provided, and LLMs are only required to make modifications to it, which tests the molecular structure knowledge of LLMs. In this case, we have crafted three subtasks for MolEdit: AddComponent, DelComponent, and SubComponent. In AddComponent, LLMs are instructed to add a specific functional group to the given molecule, and DelComponent challenges LLMs to remove a specified functional group from the provided molecule, while SubComponent is a hybrid of the previous two subtasks, requiring LLMs to first remove a designated functional group and then introduce a new one as specified to the molecule.
MolOpt challenges LLMs to not only edit molecules but also to discern whether the modification will steer the molecule towards a desired optimization target. To assess this capability, we concentrate on three pivotal properties that are vital for molecule discovery and can be effectively validated: LogP (Octanol-water partition coefficient, a metric of lipophilicity), MR (molecular refractivity, a proxy for the molar refractive index), and QED (Quantitative Estimate of Druglikeness, an assessment of drug-like characteristics). These metrics offer critical information about the potential pharmacological attributes of the molecule, which could help chemists filter molecules as viable drug candidates.
MolCustom is the most challenging task, where we have established three subtasks: AtomNum, BondNum, and FunctionalGroup. For AtomNum, LLMs are tasked with generating molecules that adhere to a specified count and type of atoms. BondNum involves the creation of molecules with a defined number and type of bonds. In FunctionalGroup, LLMs must generate a molecule that includes functional groups as specified. These subtasks may appear straightforward, yet they are deceptively challenging. They demand that LLMs have a sophisticated understanding of molecular syntax to precisely generate molecules that meet the complex criteria set forth.
Compared to previous text-based target molecule generation datasets, the task design in TOMG-Bench is rather naive, but focuses more on the details of the molecular substructure, which puts even more rigorous requirements on both the molecular understanding and generation capabilities of LLMs. Furthermore, the accurate understanding on molecule substructures and editing operations also ensures the explainability and trustworthiness of generated molecules.
3.2 Dataset Construction
Previously, molecule-related datasets and tasks have been hindered by the scarcity of human annotations. For instance, ChEBI-20  (Edwards et al., 2022), a dataset for the molecule-caption translation task, contains only 26,407 samples for training. In contrast, datasets for the text-to-image generation task in the computer vision field like MS COCO  (Chen et al., 2015) have over 1,500,000 annotated captions on more than 330,000 images. Meanwhile, the annotation of molecules demands expertise from chemists and can sometimes require wet lab experiments, which are both expensive and time-consuming. In contrast, TOMG-Bench, as an open-domain generation task, does not rely on human annotations for construction. Instead, we focus on basic molecule structural properties and molecule operations that could be validated by chemical toolboxes, which makes it easy to increase the volume of data, thereby resolving data hunger in chemical corpora.
In TOMG-Bench, three major tasks consist of nine subtasks in total. Each subtask is allocated 5,000 test samples. Figure 4 illustrates the data construction process.
@@ 3.2 Dataset Construction
For MolEdit and MolOpt, we sample molecules from a specific molecule database. We mainly consider two molecule databases: Zinc-250K (Sterling & Irwin, 2015), and PubChem  (Kim et al., 2019). Zinc-250K has 250,000 molecules, while PubChem has about 10 million molecules. Therefore, the data size of Zinc-250K is more suitable for us to build the test sets, while PubChem could be further sampled to generate the instruction tuning dataset. After sampling, we utilize RDKit (Landrum, 2013), a molecular informatics toolbox, to collect basic molecule statistics, including the structural patterns and chemical properties like LogP, MR, and QED values, which can then be integrated into our pre-defined task prompts.
For the MolCustom task, we randomly generate 5,000 instructions as requests for each subtask, requiring different numbers and collections of atoms, bonds, and functional groups. Further details are provided in the Appendix D.
3.3 Metric Design
The evaluation of the TOMG-Bench is facilitated through a set of carefully designed metrics tailored to the specific tasks within the benchmark.

### v3

# Jiatong Li [^1] footnotemark: 1    Junxian Li3    Weida Wang2,4    Yunqing Liu1    Changmeng Zheng1 Dongzhan Zhou [^2] footnotemark: 2    Xiao-yong Wei1    Qing Li1

@@ Abstract
Recently, Large Language Models (LLMs) have shown great potential in natural language-driven molecule discovery. However, existing datasets and benchmarks for molecule-text alignment are predominantly built on a one-to-one mapping, measuring LLMs’ ability to retrieve a single, pre-defined answer, rather than their creative potential to generate diverse, yet equally valid, molecular candidates. To address this critical gap, we propose Speak-to-Structure (S2-Bench), the first benchmark to evaluate LLMs in open-domain natural language-driven molecule generation. S2-Bench is specifically designed for one-to-many relationships, challenging LLMs to demonstrate genuine molecular understanding and generation capabilities. Our benchmark includes three key tasks: molecule editing (MolEdit), molecule optimization (MolOpt), and customized molecule generation (MolCustom), each probing a different aspect of molecule discovery. We also introduce OpenMolIns, a large-scale instruction tuning dataset that enables Llama-3.1-8B to surpass the most powerful LLMs like GPT-4o and Claude-3.5 on S2-Bench. Our comprehensive evaluation of 28 LLMs shifts the focus from simple pattern recall to realistic molecular design, paving the way for more capable LLMs in natural language-driven molecule discovery.
@@ 1 Introduction
Molecule discovery plays a pivotal role in various scientific research fields, from pharmaceuticals (Keiser et al., 2010) to materials science (Higuchi et al., 2023). Traditionally, this is a trial-and-error process (Ekins, 2024) that requires extensive experimentation and data analysis (Mattern & Grosser, 2023), often taking over a decade to bring a new drug candidate to market (Lee et al., 2018).
Recently, Large Language Models (LLMs) have demonstrated great potential in molecule discovery (Edwards et al., 2021; 2022) by leveraging their powerful language understanding and robust reasoning abilities (OpenAI, 2023). Molecules can be represented as SMILES strings (Weininger, 1988), which allows LLMs to process molecules as textual strings, effectively bridging the gap between molecular structures and natural language. Meanwhile, by aligning molecules with textual descriptions (Edwards et al., 2022), LLMs can assist chemists in interpreting chemical knowledge, suggesting structural modifications, and predicting compound properties (Li et al., 2024a; Zhang et al., 2024), thereby significantly streamlining the molecule discovery process.
While the integration of LLMs holds immense promise, a significant challenge lies in the limitations of current datasets and benchmarks. For example, existing datasets for aligning molecules and texts like ChEBI-20 (Edwards et al., 2022) and PubChem324K (Liu et al., 2023), are constructed based on a one-to-one mapping assumption, where each textual description is linked to a single, predefined target molecule due to the convenience of acquiring ground truth data from existing databases. However, the one-to-one mapping presents a key limitation: it does not fully align with the nature of real-world molecule discovery. In practice, multiple distinct molecules can often share the same desired properties or biological activity Petrone et al. (2012). For instance, a particular pharmaceutical effect is rarely unique to a single compound, and a material with a specific physical property, like tensile strength, can be realized through various molecular structures. This mismatch between the evaluation paradigm and the reality of molecule discovery casts doubt on whether these datasets and benchmarks genuinely assess an LLM’s capacity for creative molecular design, or if they instead inadvertently encourage models to rely on memorization and pattern-matching.
To bridge this critical gap, we introduce a novel benchmark, Speak-to-Structure (S2-Bench), the first benchmark designed to evaluate the open-domain natural language-driven molecule generation capabilities of LLMs. S2-Bench is meticulously structured into three primary tasks, as shown in Figure 1, each targeting a specific, real-world capability essential for drug and materials science:
Molecule Editing (MolEdit) focuses on testing an LLM’s ability to perform precise, localized structural modifications while preserving main structures. In drug discovery, this is analogous to Lead Optimization Barcelos et al. (2022), which edits the lead compound to create new variants. This requires a deep understanding of chemical grammar, including valency, stereochemistry, and ring systems, ensuring that edits could create valid molecules.
Molecule Optimization (MolOpt) extends the MolEdit task by evaluating a model’s ability to optimize a lead compound under specified property constraints (e.g., increasing solubility or reducing toxicity). Its open-ended design requires the model to edit and refine molecular structures in accordance with the desired properties, thereby showcasing its understanding of both structural modifications and property-related characteristics.
Customized Molecule Generation (MolCustom) requires LLMs to synthesize a novel molecule based on a natural language description of its desired structural components. Specifically, it challenges models to generate molecules with a set of quantitative and qualitative constraints, such as a specific number of atoms, bonds, or functional groups, which directly mimics the initial drug design phase where a chemist defines a new compound based on a set of precise requirements for its structure.
Figure 1: Task illustration of S2-Bench for open domain natural language-driven molecule generation. In contrast to text-based target molecule generation, multiple valid molecules may fulfill the textual requirements (right of the arrow).
To facilitate the evaluation of the above tasks, we also introduce a robust evaluation system with tailored examination process, enabling a comprehensive assessment of LLMs’ performance in open-domain natural language-driven molecule discovery. Furthermore, we propose OpenMolIns, a large-scale instruction-tuning dataset, comprising up to 120 million instruction–molecule pairs, enabling Llama3.1-8B to achieve the state-of-the-art performance among 28 powerful LLMs on S2-Bench.
To summarize, our contributions are primarily threefold:
@@
    We introduce Speak-to-Structure (S2-Bench), the first benchmark for open-domain natural language-driven molecule generation. Moving from one-to-one to one-to-many relationships, S2-Bench better aligns concepts in drug discovery and provides a novel perspective on assessing LLMs’ genuine molecular understanding and design capabilities.
@@
    We introduce OpenMolIns, an instruction-tuning dataset with up to 120 million instruction–molecule pairs, enabling Llama3.1-8B to achieve SOTA performance on S2-Bench.
@@
    We provide insightful findings based on our extensive benchmarking of 28 LLMs, revealing the limitations of existing targeted generation datasets and highlighting the potential for LLMs to transition from simple pattern recall to realistic molecular design.
2 Speak-to-Structure
In this section, we detail the design philosophy, technical composition, and statistics of Speak-to-Structure (S2-Bench), which is fundamentally structured around the core capabilities required for real-world molecular design: editing, optimizing, and customized generation.
2.1 Task Composition
S2-Bench is meticulously designed to move beyond simple pattern-matching and evaluate the genuine molecular understanding and generation capabilities of LLMs. We adopt a one-to-many paradigm, where a single language instruction can be fulfilled by multiple valid molecules. Our benchmark challenges LLMs to demonstrate a flexible understanding of molecular syntax rather than mere memorization. This design philosophy is concretely realized through three tasks of increasing complexity, each mirroring a critical phase of molecule discovery.
MolEdit assesses an LLM’s foundational molecular syntax and structural modification capabilities. In this task, a base molecule is provided, and the model is asked to perform a specific change on the molecular structure. In a real-world scenario, this task is analogous to the iterative lead optimization process in drug discovery, where chemists make small, precise changes to a candidate molecule to change its properties. We designed three subtasks to probe this capability: AddComponent and DelComponent test the model’s ability to add or remove a specific functional group, respectively, while SubComponent challenges it to perform a combined operation of both.
MolOpt evaluates an LLM’s ability to perform goal-oriented chemical reasoning. The challenge here is twofold: the LLM should not only edit a molecule but also ensure that the modification leads to a desired change in a specific property. This task directly mirrors the hit-to-lead optimization stage, where molecules are refined to achieve better pharmacological profiles. To assess this, we chose three key properties (LogP, MR, and QED), which are vital for drug discovery and can be reliably and reproducibly calculated. By requiring the model to generate a molecule that improves on a given property, we test its ability to understand the complex relationship between molecular structure and function, a capability that cannot be evaluated by tasks with a single, predetermined answer.
MolCustom serves as the ultimate test of an LLM’s creative chemical design ability. MolCustom requires the LLM to generate a novel molecule from a natural language description of its desired structural components. This task is not about modifying an existing structure but about synthesizing a new one from scratch, akin to a chemist defining the needs for a new compound at the very beginning of a project. To ensure a clear and precise evaluation of this complex task, we focus on generating molecules with specific structural features: a defined number of atoms (AtomNum), a specified number and type of bonds (BondNum), or the inclusion of particular functional groups (FunctionalGroup). While these constraints may seem simple, they are deceptively challenging, requiring the model to apply a sophisticated understanding of chemical valence rules to create a valid and novel molecule.
Compared to targeted molecule generation datasets, these tasks in S2-Bench, while seemingly straightforward in their instructions, impose a more rigorous requirement on LLMs’ ability to perform precise, open-domain molecule generation. Our benchmark fundamentally shifts the evaluation from pattern recall to practical chemical design, thereby providing a more accurate measure of an LLM’s potential in molecule discovery and helping to build more explainable and trustworthy models.
2.2 Data Construction
Previously, the development of robust datasets for text-based targeted molecule generation has been hindered by the scarcity of high-quality human annotations. For example, while text-to-image generation datasets like MS COCO (Chen et al., 2015) contain millions of annotated samples, molecule-caption datasets, such as ChEBI-20 (Edwards et al., 2022) are significantly smaller, often by orders of magnitude. This disparity arises because molecular annotation demands specialized expertise, making it both time-intensive and costly. The resulting data scarcity further poses a significant challenge for advancing LLMs in text-guided molecule discovery.
Our approach to data construction fundamentally bypasses this bottleneck. S2-Bench is designed for open-domain molecule generation tasks that do not rely on human annotations. Instead, we leverage automated chemical toolkits to programmatically construct tasks and evaluate outputs based on objective molecular properties and structural rules. This enables us to generate a virtually unlimited volume of data, effectively addressing the “data hunger” that has constrained previous research. More importantly, this programmatic approach allows us to create tasks with a one-to-many relationship, where multiple correct answers exist for a single prompt, which better aligns with the complexities of drug discovery, accommodating its inherent variability and diverse outcomes.
Specifically, we design a systematic and scalable data construction process, as illustrated in Figure 4.
@@ 2.2 Data Construction
For MolEdit and MolOpt, we sample molecules from large, publicly available chemical databases. We selected Zinc-250K (Sterling & Irwin, 2015) for building our test set due to its manageable size and diversity, while the massive PubChem (Kim et al., 2019) database (with over 10 million molecules) serves as the basis for the instruction-tuning dataset. We utilize the RDKit (Landrum, 2013) toolbox to automatically extract key molecular statistics, including structural patterns and properties like LogP, MR, and QED. These extracted features are then integrated into our pre-defined, instruction-based prompt templates, which is demonstrated in Appendix D.
For MolCustom, we move beyond existing databases and programmatically generate instructions to test a model’s ability to create novel molecules. For each of the three subtasks, we randomly generate 5,000 instructions that specify a target number and type of atoms, bonds, or functional groups.
Furthermore, for each subtask, we pre-defined a diverse prompt template pool to ensure that the LLMs are not overfit to a limited set of prompt formats. This programmatic construction not only allows for the generation of a much larger data volume but also ensures that our benchmark tests an LLM’s intrinsic chemical reasoning ability, rather than its capacity for memorizing human-annotated patterns. By moving beyond the one-to-one paradigm, S2-Bench provides a more robust and realistic evaluation, paving the way for models that can truly innovate and design novel molecules.
2.3 Evaluation
The evaluation of S2-Bench is facilitated through a set of carefully designed automated evaluation processes and metrics tailored to the unique nature of our tasks. Unlike one-to-one datasets or benchmarks that simply check for exact matches, our metrics are designed to assess an LLM’s ability to produce valid, relevant, and novel outputs within an open-ended framework.